Fifteen years from now Ravi’s age will be four times his present age. What is Ravi’s present age? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the present age of Ravi be x.
Fifteen years later, Ravi age will be x+15 years. According to the question, x+15=4x
⇒4x-x=15
⇒3x=15
⇒x=15/3
⇒x=5
Therefore, Present age of Ravi = 5 years.

Thanks